OpenStudy (anonymous):

Find the general term a_n for the geometric sequence a_1 = -3 and a_2 = 12

jimthompson5910 (jim_thompson5910):

you have the first term, and that is a_1 = -3

jimthompson5910 (jim_thompson5910):

you need the common ratio now that's just r = (second term)/(first term) r = (a_2)/(a_1) r = (12)/(-3) r = -4

jimthompson5910 (jim_thompson5910):

you now know the first term is a_1 = -3 and the common ratio is r = -4

jimthompson5910 (jim_thompson5910):

put this all together to get this nth term formula a_n = (a_1)*(r)^(n-1) a_n = -3*(-4)^(n-1) and that's all there is to it

OpenStudy (anonymous):

is it n minus 1 or is it n = -1 I am sorry I'm still a little lost this isn't my best subject

jimthompson5910 (jim_thompson5910):

it's n-1

jimthompson5910 (jim_thompson5910):

the reason it's not just 'n' is because if you start at n = 1, then you have to shift from 'n' to n - 1

jimthompson5910 (jim_thompson5910):

so the formula in a better view is \[\Large a_{n} = -3(-4)^{n-1}\]

OpenStudy (anonymous):

oh ok thats better I was getting confused now I understand it thank you

jimthompson5910 (jim_thompson5910):

Notice how if n = 1 (first term), then... \[\Large a_{n} = -3(-4)^{n-1}\] \[\Large a_{1} = -3(-4)^{1-1}\] \[\Large a_{1} = -3(-4)^{0}\] \[\Large a_{1} = -3(1)\] \[\Large a_{1} = -3\] which is the first term given above You can do the same for n = 2 to find that a_2 = 12

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

you're welcome